A switch board can handle only 4 telephone calls per minute. If the incoming calls per

minute follow poission distribution with parameter 3, find the probability that the switch

board is over taxed in any one minute?


1
Expert's answer
2021-06-15T13:51:49-0400

Let "X=" the number of telephone calls per minute: "X\sim Po(\lambda)."


"P(X=x)=\dfrac{e^{-\lambda}\cdot \lambda^x}{x!}"

Given "\lambda=3."


"P(X>4)=1-P(X=0)-P(X=1)"

"-P(X=2)-P(X=3)-P(X=4)"

"=1-\dfrac{e^{-3}\cdot 3^0}{0!}-\dfrac{e^{-3}\cdot 3^1}{1!}"

"-\dfrac{e^{-3}\cdot 3^2}{2!}-\dfrac{e^{-3}\cdot 3^3}{3!}-\dfrac{e^{-3}\cdot 3^4}{4!}"

"=1-\dfrac{e^{-3}}{8}(8+24+36+36+27)\approx0.1847"

The probability that the switch board is over taxed in any one minute is 0.1847.